# Leveraging Anthropometric Measurements to Improve Human Mesh Estimation and Ensure Consistent Body Shapes

### v3

@@ 1 Introduction
HME models are performing well on everyday data. However, in more challenging scenarios like sports, their performance is inferior to finetuned SOTA 3D HPE models. 3D HPE models only predict 3D keypoints resulting in a stickfigure pose, whereas HME models output a posed mesh including the human’s surface. Due to the lack of GT meshes, HME models cannot be trained on datasets with solely 3D keypoint annotations. The usage of synthetic data is emerging in the field, but is not applicable to challenging or specific scenarios like sports. In this paper, we propose a solution to that problem. With our A2B model and anthropometric measurements, we can now create the body shape parameters of humans needed for HME. We further apply inverse kinematics (IK) to produce the rotations that are missing in the 3D stick-figure model that is created by 3D HPE models. Together with our A2B body shape, we are able to generate human meshes that have a consistent body shape and a precise pose. The main goal of this paper is to estimate the best possible human mesh with a consistent body shape by adding the marginal overhead of measuring humans. Since our main focus is sports, this overhead is negligible, as professional athletes are commonly measured anyways. We show qualitative examples of our model and a SOTA HME model, SMPLer-X [3], in Figure 1. Our approach is generally applicable to any HME problem. We choose sports datasets to validate our proposed approach, since the poses in sports are challenging for existent HME models and athletes are measured. Their performance is currently not good enough to use them in performance assessments of athletes, which we try to change. Our contributions can be summarized as follows:
@@
  • We create and evaluate different models to convert between anthropometric measurements and SMPL-X body shape parameters. We call these models A2B. We are the first to intoduce a possibility to convert from those measurements to body shape parameters.
@@
  • With finetuned SOTA 2D and 3D HPE models [35, 10], IK, anthropometric measurements, and our A2B model, we estimate accurate human meshes with a consistent body shape. We show that this approach achieves superior results to SOTA HME models, although still evaluated on the inconsistent GT.
  • Our models and code for our approach are publicly available: https://anonymous.4open.science/r/A2B_human_mesh-FA54/
@@ 2 Related Work
Human Mesh Estimation (HME) is an active area of research. Body models like SMPL [20] and its successor SMPL-X [27] are broadly used. Their advantage is that they decouple human pose and shape. The pose parameters $\theta$ give the rotations of the joints relative to the parent joint. The shape parameters $\beta$ model the basic body shape. At first, a mesh is created with a linear mapping from $\beta$ parameters to a T-shaped pose. Next, some pose-specific shape deformations are applied and then the mesh is rotated at the joints according to the $\theta$ parameters.
@@
The first HME model that estimates SMPL-X meshes from images was introduced along SMPL-X [27]. This model, called SMPLify-X, detects 2D image features and then fits an SMPL-X model to these. To achieve that, they incorporate a pose prior trained on a large motion capture dataset and an interpenetration test. A more recent model for HME is Multi-HMR [1]. It predicts 2D heatmaps for person centers and based on that the human mesh with a human prediction head. OSX [19] is another HME model. It uses a component aware transformer that is composed of a global body encoder and local decoders for face and hands. SMPLer-X [3] is introduced as a generalist foundation model for HME, which is trained on a large amount of datasets. It mainly uses vision transformers. There is a multitude of other recent HME models, some focussing more on whole-body HME [7, 11, 23], others on multi-person HME - either with a two stage approach using a person detector and a single person human mesh estimator [6, 13, 29], or a single stage approach estimating the meshes of all persons at once [30, 34, 36].
@@
Choutas et al. [8] observed that existing HME models focus more on the body pose than the shape, although the shape is equally important for many applications. They propose SHAPY, a model that uses anthropometric and linguistic attributes to create accurate body shapes. Moreover, Sarkar et al. [32] introduce SoY, which contains specific loss functions to enhance the body shape accuracy. AnthroNet [28] propose a new body model that is learned with an end-to-end trainable pipeline. It takes anthropometric measurements as an input to learn a mesh model that accurately captures shapes of humans, but this model is different from the commonly used SMPL-X model. We use the common SMPL-X body model and decouple the estimation of the shape from the estimation of the pose. This way, we ensure the consistency of the body shape over time.
@@
In the last years, 3D HME approaches leveraged Inverse kinematics (IK) to enhance their results. HybrIK [17] transforms 3D joint coordinates to relative body-part rotations for 3D HME by using a twist-and-swing decomposition. HybrIK-X [18] further enhances HybrIK with expressive face and hands. Cha et al. [4] leverage IK to tackle the challenge of person-to-person occlusions in images with interacting persons. PLIKS [33] (Pseudo-Linear Inverse Kinematic Solver) approaches HME as a model-in-the-loop optimization problem by analytically reconstructing the human model via 2D pixel-aligned vertices in an IK-like manner.
@@ 4 From Measurements to Body Shape
Humans have been measured for centuries [9]. Tailors know exactly which measurements to take for perfectly fitting a suit or dress to the body shape of a customer. In sports, it is already common practice that professional athletes are measured for precise performance assessments. Measuring a human is easy and well understood. In contrast, the parameters of the body shape for human mesh models like SMPL-X [27] are not humanly interpretable. The $\beta$ parameters describe the principal components of the human body shape with typically around 10 to 16 values and are the result of a PCA executed on the human meshes of a training dataset while learning the SMPL-X model. Fixing all $\beta$ parameters despite one and looking at the results lets human observers get a notion of what this parameter might mean, but in total, the $\beta$ parameters and their interactions are not well interpretable. Therefore, we want to leverage the well established technique of measuring humans to create precise body shape parameters for the commonly used SMPL-X human mesh model. We call our approach to convert from 36 Anthropometric measurements to Body shape parameters A2B. Since there is no known relation between anthropometric measurements (AMs) and $\beta$ parameters, our aim is to learn this mapping. We are the first to present a method to convert from anthropometric measurements to body shape parameters. The reverse direction, B2A, is a deterministic function of the human mesh, as the AMs can be measured from the mesh.
@@ 4.2 Models
We use the same number of $\beta$ parameters for each gender as used in the AGORA dataset, meaning $11$ for male, $10$ for female, and $16$ for neutral subjects. With $36$ AMs as input values and $10-16$ output values for our A2B models, the dimensionality of the data is low. Therefore, we experiment with Support Vector Regression (SVR) and with small neural networks (NN). We split the AGORA dataset in a 80% train, 15% test and 5% validation subset. For SVR, we additionally randomly sample $10,000$ subjects for training. We use a hyperparameter search based on the validation split to determine the optimal settings, which leads us to a radial basis function kernel, an error margin of $\epsilon=0.012$ and a regularization constant of $C=3791$. For the NNs, we randomly sample new data in each iteration. The hyperparameter search for the NNs results in a model with $4$ layers, $330$ neurons per layer, tanh as activation function, and Xavier Gloriot as initialization. We use mean squared error on the model output (the $\beta$ parameters) as training loss.
@@ 4.3 Results
We train each model (NN and SVR) for each gender and with different dataset variants: We train solely on the AGORA train split, as well as on uniformly and normally distributed randomly sampled data according to the data analysis, and we further extend the range of the data as described in Section 4.1 with $\alpha_{n}=\alpha_{u}=1.5$. The results are displayed in Table 2. We evaluate the performance of our models in two ways. At first, we calculate the error of the predicted and GT $\beta$ parameters. Second, we calculate the mean deviation of the AMs of the meshes from the predicted and GT $\beta$ parameters ($A$). Therefore, this evaluation can further be seen as a kind of cycle consistency evaluation of A2B (our learned model) and B2A (the deterministic measuring function). We provide a visualization of the evaluation process in the supplementary. The anthropometric error is our main metric as these values reflect the desired body shape given as an input by the user and are further interpretable. The $\beta$ parameters are somehow arbitrary in their scale. For all genders and SVR, using an extended uniformly sampled dataset works best. For the NNs, a uniformly sampled dataset works best for male (m) and female (f) genders and an extended normally sampled dataset for the neutral (n) meshes. The results for the neutral model are worse in general, especially in the case of the NNs, which might be due to the fact that the neutral model needs to express a more diverse range of body shapes. Furthermore, the SVR achieves better results for all genders. Thus, we use these models for all datasets, without any finetuning or adaptation to specific datasets.
@@ 5.1 Improving HME Model Results
A major problem for HME based analyses is a varying basic body shape within a single video. Existing HME models output different $\beta$ parameters for each frame. Exemplarily, we show the standard deviation of the body height of one subject in Table 3. Recall that these measurements and $\beta$ parameters are based on a T-pose mesh, hence varying poses have no influence on measuring and $\beta$ parameters. Using $\beta$ parameters generated with A2B models solves this problem. The necessary 36 measurements are either measured from the human directly, or averaged from the provided GT mesh (fit3D) or IK applied to the GT poses (ASPset). We call these measurements pseudo GT and include more details in the supplementary. We choose this process to simulate real measurements which exist for most professional athletes.We combine existing HME models with the body shape estimated by our A2B models by replacing the estimated $\beta$ parameters with the ones predicted by the A2B models. We select three recent well performing models on the AGORA dataset (SMPLer-X [3], OSX [19], Multi-HMR [1]), and the first HME method developed by the SMPL-X authors, SMPLify [27]. Since SMPLer-X is trained on the official training data of fit3D, an evaluation with this model is not meaningful and we omit it here. Moreover, SMPLify-X is not SOTA any more and achieved the worst results for ASPset. Therefore we omit it, too.
@@
Although it is not our main goal, we further evaluate the capabilities of a fixed body shape without available GT measurements to ensure consistent body shapes in the case that no measurments are available. The simplest approach is to use the median of the $\beta$ parameters across all frames of the respective model. However, the $\beta$ parameters have no real meaning. Therefore, we compare this approach to taking the median of the anthropometric measurments of the generated meshes and then converting them to $\beta$ parameters via the A2B models. Results are displayed in Table 4. For SMPLer-X and OSX, using the median $\beta$ parameters lead to equal or even worse results on ASPset. Regarding ASPset, using our A2B models increases the performance of all models slightly. Switching from the neutral output that these models all have to a gendered model works best in most of these cases, but the neutral A2B models also lead to a marginal improvement. Regarding fit3D, using the median $\beta$ parameters already enhances the MPJPE and MVE results. Using $\beta$ parameters from an A2B model leads to the same improvement for both metrics, OSX achieves the best results with SVR and the neutral model, Multi-HMR with NN and the neutral model.
@@
Anthropometric measurements can further be used to easily convert between neutral and gendered (male or female) models. In contrast, $\beta$ parameters are not transferable between models of different genders. Therefore, until now, the conversion could only be achieved by minimizing the MVE between meshes of different genders in an iterative process. We can now use the B2A fuction to obtain measurements for a mesh of one gender and apply the A2B model of the other gender to these anthropometric measurements in order to get the corresponding $\beta$ parameters for this gender.
@@ 5.2 HME with Sequence Based 3D HPE and A2B
All evaluated HME models are working image-wise. In contrast, SOTA 3D HPE models take a long sequence of 2D poses as an input, which helps to capture movements precisely. The models are called uplifting models, since they lift 2D pose sequences to 3D pose sequences. We use the efficent SOTA 3D HPE model uplift and upsample (UU) [10] to estimate the 3D poses on videos. To estimate the required 2D poses from the video frames, we use ViTPose [35], a SOTA 2D pose estimation model. It is important to note that UU operates on pose sequences instead of single frames like the HME models in Section 5.1 and can leverage the information of neighboring frames to estimate a more sophisticated pose. Since we have GT 3D joints available, we can finetune the models (ViTpose for 2D HPE and UU for 3D HPE) on our data. This is also necessary to adapt the model to the dataset specific joint definitions since many 3D HPE models like UU are pretrained on datasets like Human3.6m [16], but those joint definitions do not match ASPset nor fit3D. We finetune both 2D and 3D HPE models on the training subsets. On the test subsets, UU achieves an MPJPE of 63.85 mm on ASPset and an MPJPE of 29.60 mm on fit3D, which is better than the best existing HME model for both datasets (see Section 5.1). However, UU only outputs 3D joints, no meshes. Moreover, a stick-figure 3D pose is not sufficient to model the pose parameters $\theta$ of the SMPL-X mesh, since some rotations are missing. Hence, it is impossible to calculate the necessary rotation parameters directly from the UU result.
@@ 5.2.1 Inverse Kinematics for Full Pose Estimation
Therefore, we use the well established approach of inverse kinematics (IK) with a pose prior to obtain the missing rotations by fitting an SMPL-X mesh to the 3D joint locations estimated by UU. IK outputs the best SMPL-X parameters ($\beta$ and $\theta$) that fit the mesh to the given 3D joint locations. Details can be found in the supplementary.
@@ 5.2.2 Experiments and Results
Our main approach is shown in the second rows in Table 5. We evaluate the results of IK applied to the UU joint locations with original, median, and pseudo GT based A2B $\beta$ parameters. The real-world scenario corresponds to the following. GT measurements can be measured from the athlete directly and the 3D pose can be estimated with UU and IK. The $\beta$ parameters are estimated with the A2B models. A visualization of this pipeline can be found in Figure 3. We include the best result(s) from existing HME models in the respective last rows for comparison and provide qualitative results in Figure 1. Our approach outperforms the best existing HME model for both datasets by a large margin. But we need to mention that our approach is not directly comparable to the original results of the HME models, since our approach needs the additional information of measurements, but they already exist in our scenario. However, our approach still outperforms the existing HME models even when they use the same measurments and A2B results as our model does. Further, our model provides results for all frames, which is not the case for the other HME models.
@@
Next, we replace the inconsistent $\beta$ parameters with the results from our A2B models. This is especially helpful for our approach since IK produces body shapes with high inconsistencies, as shown by the larger standard deviation of the body height compared to other HME models. For ASPSet, using pseudo GT AMs results in a large improvement of over 12 mm. Remarkably, this result surpasses even the original UU result by 8 mm. It seems that incorporating a clearly defined mesh helps to fix some typical errors of UU and enhance its result in case of ASPset. In general, the error on fit3D is much lower for UU based approaches. The reason might be that it consists of much more data, such that we can finetune UU for a longer time. Further, the videos are recorded in a lab in comparison to the in-the-wild videos of ASPset. The lab environment is very similar to the Human3.6m dataset [16], which serves as a training dataset for most recent HME models. Therefore, the results of ASPset are more relevant for future applications of our approach, where we assume only a few available 3D annotations and in-the-wild recordings. For fit3D, applying the A2B body shapes from pseudo GT AMs leads to a slight decrease in performance of 0.2 mm. Inconsistent shapes in the GT (see Section 3) are likely to cause this behavior. Still, our approach using a 3D HPE model and IK outperforms all existing HME models, no matter if the original inconsistent or the consistent body shapes from A2B are used.
@@
As described in Section 5.1, we further evaluate the capabilities of a consistent shape without available GT AMs. The naive approach is to use the median of the estimated inconsistent $\beta$ parameters (Tab. 5, column median). Another approach is to use the meshes created by IK applied to the UU results, compute the AMs with B2A, calculate the median AMs and convert them to $\beta$ parameters via the A2B models. Results are displayed in Table 5, row three. For ASPset, using fixed body shape parameters from A2B models based on the measurements from UU results achieves a slightly better score than the results with inconsistent body shapes. For fit3d, the MPJPE increases by 0.9 mm, but the A2B model results are a slighly better alternative for consistent body shapes compared to the median $\beta$ parameters.
Further, our approach can be used to generate pseudo GT meshes for datasets with only 3D keypoint annotations. We can use these pseudo GT meshes to finetune HME models and increase their performance regarding the estimated keypoints for the specific dataset. However, these results are still worse than the results of our approach. We present the results in the supplementary.

### v4

@@ 1 Introduction
HME models are performing well on everyday data. However, in more challenging scenarios like sports, their performance is inferior to fine-tuned SOTA 3D HPE models. 3D HPE models only predict 3D keypoints resulting in a stick-figure pose, whereas HME models output a posed mesh including the human’s surface. Due to the lack of GT meshes, HME models cannot be trained on datasets with solely 3D keypoint annotations. The usage of synthetic data is emerging in the field, but is not applicable to challenging or specific scenarios like sports. In this paper, we propose a solution to that problem. With our A2B model and anthropometric measurements, we can now create the body shape parameters of humans needed for HME. We further apply inverse kinematics (IK) to produce the rotations that are missing in the 3D stick-figure model that is created by 3D HPE models. Together with our A2B body shape, we are able to generate human meshes that have a consistent body shape and a precise pose. The main goal of this paper is to estimate the best possible human mesh with a consistent body shape by adding the marginal overhead of measuring humans. Since our main focus is sports, this overhead is negligible, as professional athletes are commonly measured anyway. We show qualitative examples of our model and a SOTA HME model, SMPLer-X [3], in Figure 1. Our approach is generally applicable to any HME problem. We choose sports datasets to validate our proposed approach, since the poses in sports are challenging for existent HME models and athletes are measured. Their performance is currently not good enough to use them in performance assessments of athletes, which we try to change. Our contributions can be summarized as follows:
@@
  • We create and evaluate different models to convert between anthropometric measurements and SMPL-X body shape parameters for all genders. We call them A2B.
@@
  • With fine-tuned SOTA 2D and 3D HPE models [37, 10], IK, anthropometric measurements, and our A2B model, we estimate accurate human meshes with a consistent body shape. We show that this approach achieves superior results to SOTA HME models, although still evaluated on the inconsistent GT.
  • Our models and code for our approach are publicly available: https://github.com/kaulquappe23/a2b_human_mesh
@@ 2 Related Work
Human Mesh Estimation (HME) is an active area of research. Body models like SMPL [20] and its successor SMPL-X [27] are broadly used. Their advantage is that they decouple human pose and shape. The pose parameters $\theta$ give the rotations of the joints relative to the parent joint. The shape parameters $\beta$ model the basic body shape. At first, a mesh is created with a linear mapping from $\beta$ parameters to a T-shaped pose. Next, some pose-specific shape deformations are applied, and then the mesh is rotated at the joints according to the $\theta$ parameters.
@@
The first HME model that estimates SMPL-X meshes from images, SMPLify-X, was introduced along SMPL-X [27]. It detects 2D image features and then fits an SMPL-X model to these. To achieve that, they incorporate a pose prior trained on a large motion capture dataset and an interpenetration test. A more recent model for HME is Multi-HMR [1]. It predicts 2D heatmaps for person centers and based on that the human mesh with a human prediction head. OSX [19] is a HME model using a component aware Transformer that is composed of a global body encoder and local decoders for face and hands. SMPLer-X [3] is introduced as a generalist foundation model for HME trained on a large amount of datasets mainly using vision transformers. There are many other HME models, some focussing more on whole-body HME [7, 11, 23], others on multi-person HME - either with a two stage approach using a person detector and a single person human mesh estimator [6, 13, 29], or a single stage approach estimating the meshes of all persons at once [30, 36, 38].
@@
Choutas et al. [8] observed that existing HME models focus more on the body pose than the shape, although the shape is equally important for many applications. They propose SHAPY, a model that uses anthropometric and linguistic attributes to create accurate body shapes. Moreover, Sarkar et al. [32] introduce SoY, which contains specific loss functions to enhance the body shape accuracy. AnthroNet [28] propose a new body model that is learned with an end-to-end trainable pipeline. It takes anthropometric measurements as an input to learn a mesh model that accurately captures shapes of humans, but this model is different from the commonly used SMPL-X model. We use the common SMPL-X body model and decouple the estimation of the shape from the estimation of the pose. Sengupta et al. [34] estimate anthropometric measurements from images and use a linear layer to convert them to body shape parameters. However, their model operates on single images and their measurement to shape conversion is different from ours and only available for a single gender. We further ensure the consistency of the body shape over time.
@@
In the last years, 3D HME approaches leveraged Inverse kinematics (IK) to enhance their results. HybrIK [17] transforms 3D joint coordinates to relative body-part rotations for 3D HME by using a twist-and-swing decomposition. HybrIK-X [18] further enhances HybrIK with expressive face and hands. Cha et al. [4] leverage IK to tackle the challenge of person-to-person occlusions in images with interacting persons. PLIKS [35] (Pseudo-Linear Inverse Kinematic Solver) approaches HME by analytically reconstructing the human model via 2D pixel-aligned vertices in an IK-like manner.
@@ 4 From Measurements to Body Shape
Humans have been measured for centuries [9]. Tailors know exactly which measurements to take for perfectly fitting a suit or dress to the body shape of a customer. In sports, it is already common practice that professional athletes are measured for precise performance assessments. Measuring a human is easy and well understood. In contrast, the parameters of the body shape for human mesh models like SMPL-X [27] are not humanly interpretable. The $\beta$ parameters describe the principal components of the human body shape with typically around 10 to 16 values and are the result of a PCA executed on the human meshes of a training dataset while learning the SMPL-X model. Fixing all $\beta$ parameters despite one and looking at the results lets human observers get a notion of what this parameter might mean, but in total, the $\beta$ parameters and their interactions are not well interpretable. Therefore, we want to leverage the well established technique of measuring humans to create precise body shape parameters for the commonly used SMPL-X human mesh model. We call our approach to convert from 36 Anthropometric measurements to Body shape parameters A2B. Since there is no known relation between anthropometric measurements (AMs) and $\beta$ parameters, our aim is to learn this mapping. The reverse direction, B2A, is a deterministic function of the human mesh, as the AMs can be measured from the mesh.
@@ 4.2 Models
We use the same number of $\beta$ parameters for each gender as used in the AGORA dataset, meaning $11$ for male, $10$ for female, and $16$ for neutral subjects. With $36$ AMs as input values and $10-16$ output values for our A2B models, the dimensionality of the data is low. Therefore, we experiment with Support Vector Regression (SVR) and with small neural networks (NN). We split the AGORA dataset in an 80% train, 15% test and 5% validation subset. For SVR, we additionally randomly sample $10,000$ subjects for training. We use a hyperparameter search based on the validation split to determine the optimal settings, which leads us to a radial basis function kernel, an error margin of $\epsilon=0.012$ and a regularization constant of $C=3791$. For the NNs, we randomly sample new data in each iteration. The hyperparameter search for the NNs results in a model with $4$ layers, $330$ neurons per layer, tanh as activation function, and Xavier Glorot as initialization. We use mean squared error on the model output (the $\beta$ parameters) as training loss.
@@ 4.3 Results
We train each model (NN and SVR) for each gender and with different dataset variants: We train solely on the AGORA train split, as well as on uniformly and normally distributed randomly sampled data according to the data analysis, and we further extend the range of the data as described in Section 4.1 with $\alpha_{n}=\alpha_{u}=1.5$. The results are displayed in Table 2. We evaluate the performance of our models in two ways. At first, we calculate the error of the predicted and GT $\beta$ parameters. Second, we calculate the mean deviation of the AMs of the meshes from the predicted and GT $\beta$ parameters ($A$). Therefore, this evaluation can further be seen as a kind of cycle consistency evaluation of A2B (our learned model) and B2A (the deterministic measuring function). We provide a visualization of the evaluation process and evaluation results of the real-world SSP-3D dataset [33] in the supplementary. The anthropometric error is our main metric as these values reflect the desired body shape given as an input by the user and are further interpretable. The $\beta$ parameters are somehow arbitrary in their scale. For all genders and SVR, using an extended uniformly sampled dataset works best. For the NNs, a uniformly sampled dataset works best for male (m) and female (f) genders and an extended normally sampled dataset for the neutral (n) meshes. The results for the neutral model are worse in general, especially in the case of the NNs, which might be due to the fact that the neutral model needs to express a more diverse range of body shapes. Furthermore, the SVR achieves better results for all genders. Thus, we use these models for all datasets, without any fine-tuning or adaptation to specific datasets.
@@ 5.1 Improving HME Model Results
A major problem for HME based analyses is a varying basic body shape within a single video. Existing HME models output different $\beta$ parameters for each frame. Exemplarily, we show the standard deviation of the body height of one subject in Table 3. Recall that these measurements and $\beta$ parameters are based on a T-pose mesh, hence varying poses have no influence on measuring and $\beta$ parameters. Using $\beta$ parameters generated with A2B models solves this problem. The necessary 36 measurements are either measured from the human directly, or averaged from the provided GT mesh (fit3D) or IK applied to the GT poses (ASPset). We call these measurements pseudo GT and include more details in the supplementary. We choose this process to simulate real measurements which exist for most professional athletes. We combine existing HME models with the body shape estimated by our A2B models by replacing the estimated $\beta$ parameters with the ones predicted by the A2B models. We select three recent well performing models on the AGORA dataset (SMPLer-X [3], OSX [19], Multi-HMR [1]), and the first HME method developed by the SMPL-X authors, SMPLify [27]. Since SMPLer-X is trained on the official training data of fit3D, an evaluation with this model is not meaningful, and we omit it here. Moreover, SMPLify-X is not SOTA anymore and achieved the worst results for ASPset. Therefore we omit it, too.
@@
Although it is not our main goal, we further evaluate the capabilities of a fixed body shape without available GT measurements to ensure consistent body shapes in the case that no measurements are available. The simplest approach is to use the median of the $\beta$ parameters across all frames of the respective model. However, the $\beta$ parameters have no real meaning. Therefore, we compare this approach to taking the median of the anthropometric measurements of the generated meshes and then converting them to $\beta$ parameters via the A2B models. Results are displayed in Table 4. For SMPLer-X and OSX, using the median $\beta$ parameters lead to equal or even worse results on ASPset. Regarding ASPset, using our A2B models increases the performance of all models slightly. Switching from the neutral output that these models all have to a gendered model works best in most of these cases, but the neutral A2B models also lead to a marginal improvement. Regarding fit3D, using the median $\beta$ parameters already enhances the MPJPE and MVE results. Using $\beta$ parameters from an A2B model leads to the same improvement for both metrics, OSX achieves the best results with SVR and the neutral model, Multi-HMR with NN and the neutral model.
@@
Anthropometric measurements can further be used to easily convert between neutral and gendered (male or female) models. In contrast, $\beta$ parameters are not transferable between models of different genders. Therefore, until now, the conversion could only be achieved by minimizing the MVE between meshes of different genders in an iterative process. We can now use the B2A function to obtain measurements for a mesh of one gender and apply the A2B model of the other gender to these anthropometric measurements in order to get the corresponding $\beta$ parameters for this gender.
@@ 5.2 HME with Sequence Based 3D HPE and A2B
All evaluated HME models are working image-wise. In contrast, SOTA 3D HPE models take a long sequence of 2D poses as an input, which helps to capture movements precisely. The models are called uplifting models, since they lift 2D pose sequences to 3D pose sequences. We use the efficient SOTA 3D HPE model uplift and upsample (UU) [10] to estimate the 3D poses on videos. To estimate the required 2D poses from the video frames, we use ViTPose [37], a SOTA 2D pose estimation model. It is important to note that UU operates on pose sequences instead of single frames like the HME models in Section 5.1 and can leverage the information of neighboring frames to estimate a more sophisticated pose. Since we have GT 3D joints available, we can fine-tune the models (ViTpose for 2D HPE and UU for 3D HPE) on our data. This is also necessary to adapt the model to the dataset specific joint definitions since many 3D HPE models like UU are pretrained on datasets like Human3.6M [16], but those joint definitions do not match ASPset nor fit3D. We fine-tune both 2D and 3D HPE models on the training subsets. On the test subsets, UU achieves an MPJPE of 63.85 mm on ASPset and an MPJPE of 29.60 mm on fit3D, which is better than the best existing HME model for both datasets (see Section 5.1). However, UU only outputs 3D joints, no meshes. Moreover, a stick-figure 3D pose is not sufficient to model the pose parameters $\theta$ of the SMPL-X mesh, since some rotations are missing. Hence, it is impossible to calculate the necessary rotation parameters directly from the UU result.
@@ 5.2.1 Inverse Kinematics for Full Pose Estimation
Because we need the rotation parameters, we use the well established approach of inverse kinematics (IK) with a pose prior to obtain the missing rotations by fitting an SMPL-X mesh to the 3D joint locations estimated by UU. IK outputs the best SMPL-X parameters ($\beta$ and $\theta$) that fit the mesh to the given 3D joint locations. Details can be found in the supplementary.
@@ 5.2.2 Experiments and Results
Our main approach is shown in the second rows in Table 5. We evaluate the results of IK applied to the UU joint locations with original, median, and pseudo GT based A2B $\beta$ parameters. The real-world scenario corresponds to the following. GT measurements can be measured from the athlete directly and the 3D pose can be estimated with UU and IK. The $\beta$ parameters are estimated with the A2B models. A visualization of this pipeline can be found in Figure 3. We include the best result(s) from existing HME models in the respective last rows for comparison and provide qualitative results in Figure 1. Our approach outperforms the best existing HME model for both datasets by a large margin. But we need to mention that our approach is not directly comparable to the original results of the HME models, since our approach needs the additional information of measurements, but they already exist in our scenario. However, our approach still outperforms the existing HME models even when they use the same measurements and A2B results as our model does. Further, our model provides results for all frames, which is not the case for the other HME models.
@@
Next, we replace the inconsistent $\beta$ parameters with the results from our A2B models. This is especially helpful for our approach since IK produces body shapes with high inconsistencies, as shown by the larger standard deviation of the body height compared to other HME models. For ASPset, using pseudo GT AMs results in a large improvement of over 12 mm. Remarkably, this result surpasses even the original UU result by 8 mm. It seems that incorporating a clearly defined mesh helps to fix some typical errors of UU and enhance its result in case of ASPset. In general, the error on fit3D is much lower for UU based approaches. The reason might be that it consists of much more data, such that we can fine-tune UU for a longer time. Further, the videos are recorded in a lab in comparison to the in-the-wild videos of ASPset. The lab environment is very similar to the Human3.6M dataset [16], which serves as a training dataset for most recent HME models. Therefore, the results of ASPset are more relevant for future applications of our approach, where we assume only a few available 3D annotations and in-the-wild recordings. For fit3D, applying the A2B body shapes from pseudo GT AMs leads to a slight decrease in performance of 0.2 mm. Inconsistent shapes in the GT (see Section 3) are likely to cause this behavior. Still, our approach using a 3D HPE model and IK outperforms all existing HME models, no matter if the original inconsistent or the consistent body shapes from A2B are used.
@@
As described in Section 5.1, we further evaluate the capabilities of a consistent shape without available GT AMs. The naive approach is to use the median of the estimated inconsistent $\beta$ parameters (Tab. 5, column median). Another approach is to use the meshes created by IK applied to the UU results, compute the AMs with B2A, calculate the median AMs and convert them to $\beta$ parameters via the A2B models. Results are displayed in Table 5, row three. For ASPset, using fixed body shape parameters from A2B models based on the measurements from UU results achieves a slightly better score than the results with inconsistent body shapes. For fit3D, the MPJPE increases by 0.9 mm, but the A2B model results are a slightly better alternative for consistent body shapes compared to the median $\beta$ parameters.
Further, our approach can be used to generate pseudo GT meshes for datasets with only 3D keypoint annotations. We can use these pseudo GT meshes to fine-tune HME models and increase their performance regarding the estimated keypoints for the specific dataset. However, these results are still worse than the results of our approach. We present the results in the supplementary.